onsider the following two person static game where Player 1 is the row player and Player 2 is the column player

C D E
A 1,1 0,2 2,0
B 0,0 1,-1 -1,3
a.

There is an equilibrium where Player 1 plays A with probability 3/4.

b.

There is an equilibrium where Player 1 plays A with probability 2/3.

c.

There is an equilibrium where Player 1 plays A with probability 1/2.

d.

There is no mixed strategy Nash equilibrium.
Answer #1

C

D

E

A

1 1

0 2

2 0

B

0 0

1 -1

-1 3

We assign probability p and 1-p to strategy A and B respectively of Player 1. Similarly, we assign probability of q1 , q2 and 1-q1-q2 to strategy C , D and E respectively of Player 2.

To derive the mixed strategy nash equilibrium: we need to calculate the expected pay off of the players.

Player 1 chooses probability p such that player 2 is indifferent between C and D

Thus,

p. (1) + 1-p (0) = p(2)+1-p(-1)

p=2p-1+p

p-3p = -1

p=1/2

Similarly, Player 1 chooses probability p such that player 2 is indifferent between D and E

p(2)+1-p(-1) = p(0) + 1-p (3)

2p-1+p= 3-3p

3p-1= 3-3p

6p = 4

p=2/3

This is impossible. Therefore we donot have mixed strategy naash equilibrium for this game